OpenStudy (anonymous):

Find a simple transformation for all points except those indicated: f(x)=(radical(x))-1/radical(x) / 1-(1/radical(x)); g(x)=radicalx; x=0,1

OpenStudy (anonymous):

\[f(x)=\frac{\frac{\sqrt{x}-1}{\sqrt{x}}}{1-\frac{1}{\sqrt{x}}}\]

OpenStudy (anonymous):

like that???

OpenStudy (anonymous):

ah no

OpenStudy (anonymous):

radcialx then minus 1/radicalx. denominator is good though

OpenStudy (anonymous):

\[f(x)=\frac{\sqrt{x}-\frac{1}{\sqrt{x}}}{1-\frac{1}{\sqrt{x}}}\]

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

ok lets multiply top and bottom by \[\sqrt{x}\]

OpenStudy (anonymous):

\[f(x)=\frac{x-1}{\sqrt{x}-1}\]

OpenStudy (anonymous):

this is good at 0 because you get \[f(0)=1\]

OpenStudy (anonymous):

but it is still not good at 1 because you get 0/0 so now lets rationalize the denominator by multiplying top and bottom by \[\sqrt{x}+1\]

OpenStudy (anonymous):

quick question howd u know to rationalize the function with radical x+1?

OpenStudy (anonymous):

\[\frac{x-1}{\sqrt{x}-1}\times \frac{\sqrt{x}+1}{\sqrt{x}+1}\] \[\frac{(x-1)(\sqrt{x}+1)}{x-1}=\sqrt{x}+1\]

OpenStudy (anonymous):

how do i know it? good question. but it is the same way you always rationalize a denominator or numerator. multiply by the "conjugate"

OpenStudy (anonymous):

it works because \[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

so how does that equation simplify into 1+radicalx?

OpenStudy (anonymous):

ohw wait i think i see it!

OpenStudy (anonymous):

cancel the x - 1 top and bottom

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw